OpenStudy (anonymous):

Your starship, the Aimless Wanderer, lands on the mysterious planet Mongo. As chief scientist-engineer, you make the following measurements: a 2.50-kg stone thrown upward from the ground at 15.0 m/s returns to the ground in 6.00 s; the circumference of Mongo at the equator is 4.00×105 km; there is no appreciable atmosphere, and the planet's mass is 3.04*10^26 kg. If the Aimless Wanderer goes into a circular orbit 3.00×104 km above the surface of Mongo, how many hours will it take the ship to complete one orbit?

OpenStudy (mani_jha):

\[Time=Distance(2pir)/Speed(\sqrt{GM/r)}\] r=Radius of Planet+3X10^4

OpenStudy (mani_jha):

Clear?

OpenStudy (anonymous):

first u got to find the velocity with which the man orbits the planet centripetal force is provided by the graviational force of attraction GmM/r^2 so mv^2/r=GmM/r^2 find expression for velocity next speed=distance/time so time=distance/speed i know speed wat abt the distance? man goes in a circular orbit hence the orbits radius is R perimeter=2*pi*R  note the diff btw R and r(radius of planet)